# Discovery of Volatile Gas in the Giant Impact Disk around the 150-Myr old HD 23514

### v1

@@ 3.4 Is CO2 present in the Spitzer/IRS spectrum?
Inspired by the discovery of CO2 gas, we went back to examine whether a similar gas emission was also present in the Spitzer IRS spectrum taken 15 years earlier. We retrieved the calibrated spectrum from the Combined Atlas of Sources with Spitzer IRS spectra (CASSIS)222https://cassis.sirtf.com/atlas/ (Lebouteiller et al., 2011) site. We used the optimal extraction product best for point sources and adopted the uncertainty including both statistical (rms) and systematic errors. The IRS data shown in Figure 3 do show a bump near 15 µm, which was not found in other similar debris systems (for details see Appendix C.4), suggesting that the bump is likely astrophysical in nature. We then determined the general continuum by fitting the CO2 complex region using a polynomial function and forcing the flux between 13.9 and 15.8 µm to be relatively flat for a conservative approach as shown in Figure 3. We then constructed a low-resolution model spectrum by (1) adding the JWST CO2 model (from Section 3.3) to the continuum, and (2) re-sampling the wavelength grid to match the Spitzer IRS spectral resolution333see the IRS Instrument Handbook for details. As shown in Figure 3, the low-resolution model spectrum resembles the observed one well if the same level of the CO2 emission detected by JWST was also present. In this case, the strong CO2 Q-branch was detected at 4.3 statistically (2.5 if using the combined error) while the others were buried in the noise. In summary, the same level of the hot CO2 emission is also present in the Spitzer data taken 15 years earlier and remains unchanged between the two epochs.
@@ 4.1 Colocation of silica and hot gas
The HD 23514 system has been observed with the near-IR adaptive optics imaging at the 10 m Keck II telescope (Rodriguez et al., 2012) and with ALMA at 1.3 mm using a beam size of 15 (Sullivan et al., 2022), and neither resolves nor detects the disk. Despite the high degeneracy, the disk extent can be estimated using spectral energy distribution (SED) models because the bulk of disk emission is sensitive to its temperature (i.e., location for a given dust composition). Vican et al. (2016) characterized the HD 23514 system as a two-temperature disk with a hot component having dust temperatures () of 1082 K and a warm component with of 168 K, roughly consistent with our estimated temperatures for the featureless dust continuum (750 + 200 K; Appendix A). Given the stellar properties, these temperatures correspond to radial locations of 0.1 and 4.8 au, respectively, for blackbody emitters. For imperfect emitters like dust grains, the probable location is likely to be larger (Figure 9). Under the optically thin condition, the contrast/ratio between the prominent 10 and 20 µm dust features is a good proxy for dust temperature as that the larger the ratio the hotter the dust is due to the Planck emission (Figure 9). In other words, the bulk of the dust emission that gives rise to the silica feature most likely comes from the hot rather than the warm component.
We explored SED models, aiming to constrain the location of the silica dust (for details see Appendix D). Although such models might be degenerate, they provide diagnostic constraints under simplified assumptions. We found that the minimum grain size in the disk needs to be 0.1–0.5 µm in order to reproduce the 10/20 µm feature strengths, which is smaller than the typical blowout size, 1 µm for HD 23514 (Appendix D.2). Our SED models suggest that the disk is compact (0.1–3 au), and the highly refractory silica grains are mostly confined within the sub-au region, where the temperature is close to their sublimation temperature, in order to reproduce the prominent 9 µm feature and 3–5 µm flux. The system’s infrared fractional luminosity is 1.76 with 70% coming from the featureless dust continuum. The total dust mass is g (1.67) and the feature-producing grains account for 5% of the dust mass with only 1% coming from tiny silica dust. Based on the model parameters for the gas (Section C.1) and the silica dust, they are likely co-located very close to the star at sub-au distance.
@@ 4.2 Origins of the Stable Silica Dust
The origin of the large amount of sub-µm grains in debris disks has been long debated in the literature. The typical radiation pressure blowout size is 1 µm around HD 23514 except for highly porous grains (Arnold et al., 2019), i.e., the grains that give rise to the prominent mid-IR feature have very short lifetimes (less than a year, Appendix D.3). This implies a very high mass loss rate (51022 g yr-1) if there were no optical depth effects either to prevent them leaving the system by shielding or to replenish them fast enough (e.g., Thebault & Kral 2019). If this high rate was sustained through collisional cascades over the age of the system (150 Myr), HD 23514 would need to be born with a disk that is 35 times of the minimum mass of solar nebula (MMSN), which is not impossible but unlikely. Instead, the large amount of small grains is more likely to be created by transient events where the high rate is only applicable for a limited recent time. The maximum dust production rates in solar system comets can reach 1013 g yr-1 when they are very close to the Sun. It is then unlikely to sustain the high dust production/loss rate by inward scattered comets if these grains are continuously being ejected by radiation pressure.
@@
One easy way to address this dilemma is that all or some portion of the small grains are not subject to radiation pressure blowout, which can be achieved by certain grain properties or being trapped by other means (Rieke et al., 2016; Lebreton et al., 2013; Pearce et al., 2020). The prominent 9 µm feature is thought to come from obsidian or annealed/fused quartz grains (Fujiwara et al., 2012) whose properties make them transparent to stellar photons, therefore not subject to blowout, when they are tiny enough (Artymowicz, 1988). Giant impacts provide perfect conditions for the tiny grain formation. These tiny grains could have been generated either by direct condensation from the impact-produced vapor (Johnson & Melosh, 2012) or by hyper-velocity grinding within optically thick clouds of debris produced by violent collisions of large bodies (Johnson et al., 2012). In the former case, these silica grains were born tiny; while in the later case, tiny grains were formed through collisional cascades and the optical depth effects within the formation region allow grains that are much smaller than the blowout size to form. As a result, the solid-state feature rising from the tiny silica dust would remain stable until they slowly drift inward due to the Poynting-Robertson drag (Su et al., 2020). The feature would gradually weaken due to dust sublimation once the grains reach the sublimation temperature under timescales of a few hundred years in HD 23514.
@@
In addition, the giant impact ejecta would also form a population of planetesimals, maintaining a high level of dustiness and creating temporal variation (Watt et al., 2024) as probed by 3-5 µm photometry. The observed variability levels require a minimal corresponding area change in the dust cross section to be on the order of (1–2.5)10-3 au2 (for details see Appendix B). This dust cross section is similar to the emitting area of hot gas, pointing a potential link between the dust and molecular gas production in the sub-au region. Assuming the molecular gas and small dust share the same sub-au region, the estimated gas and dust masses suggest that the small grains could have low (1) Stokes number and coupled with the gas (highly depending on the estimated CO mass, see Appendix C.1). The presence of molecular gas would increase the dust-clearing timescale due to the combination of radiation pressure and gas drag (Takeuchi & Artymowicz, 2001; Kenyon et al., 2016). Exploring grain trapping is beyond the scope of this paper, but it is likely to be an important component to a comprehensive model for the HD 23514 system, particularly in light of the new discovery of gas in the system.
@@ 4.3 Origins of the Volatile Material
In-situ Formation In the context of the giant impact hypothesis, molecular gas could be produced by outgassing processes if one or both the impacting bodies are volatile-rich. Earth’s early atmosphere was thought to form by retention of these outgassing products (Abe & Matsui, 1985). Impact experiments and model calculations both show that H2O and CO2 are the major gases produced by outgassing of carbonaceous chondritic material over a wide range of temperatures and pressures (e.g., Lange & Ahrens 1982; Schaefer & Fegley 2010; Thompson et al. 2021, and references within). Furthermore, shock-recovery experiments targeting a mixture of olivine, iron and water also reveal preferential formation of ultrafine (tens of nanometers) smoke particles of silicates and metal oxides, facilitated by the role of supercritical water (Furukawa et al., 2007). These tiny grains might be ubiquitous in post ocean impact events and provide effective UV shielding for the volatiles. Alternatively, the volatiles could also be released from (1) an atmospheric stripping event as has been suggested for HD 172555 (Schneiderman et al., 2021), or (2) the large remnants of a giant impact ejecta as long as they are large enough to retain a significant fraction of volatiles. It is then interesting to note that the change in the dust cross section (derived in Appendix B) is on the same order as the emitting area of the molecular gas, indicating that collisional activities in the aftermath of giant impact events could be tied to the formation/protection of molecular gas in the inner region.
@@
Delivery from Outer Region This delivery scenario has been widely accepted in both our solar system (comets) and other debris systems (exocomets) – scattered icy bodies from the outer region releasing volatiles and dust grains as they warm up on their inward migration. In this case, the detected molecular gas might have no direct relationship with the giant-impact-produced small grains observed in the system. For HD 23514, the inward delivery rate might be much lower than that of the debris systems that host Kuiper-belt analogs such as Pic (Beust et al., 1996; Matrà et al., 2019) and Corvi (Marino et al., 2017) because it has a limited volume of cold reservoir for storing icy bodies (lack of a massive Kuiper-belt analog, see Appendix D). This scenario also needs a perturber either internally or externally to efficiently scatter icy bodies into the inner region. Ground-based high contrast imaging found no other sources within 10″ around HD 23514 except for the M7 brown dwarf companion (Rodriguez et al., 2012) 264 away, while a recent RV study rules out massive giant planets within 0.1 au (Takarada et al., 2020). External companions under some favorable conditions can excite the planetesimal population via the eccentric Kozai–Lidov mechanism, leading to inner dust and gas production (Naoz, 2016; Young & Wyatt, 2024). Without detailed orbital parameters, it is not clear whether the M-dwarf companion at 360 au away is capable of this role. Other challenges in the delivery scenario are: (1) evaporation of icy bodies is expected to occur at a range of distances, which is at odds with the gas concentration within the sub-au region, and (2) the comet delivery rate is expected to be variable based on the solar system observations, which somehow contradicts the tentative CO2 detection in the Spitzer data. The second challenge might be accommodated if scattering of many comets at a high efficiency. A future investigation is needed to further explore the delivery model; nonetheless, it remains a possible origin for the molecular gas.
@@ 5 Conclusion
We have presented a JWST MIRI/MRS 5–28 µm spectrum of HD 23514, an F5V star in the 150 Myr-old Pleiades cluster, hosting one of the giant-impact disks discovered by Spitzer with unique mineralogy dominated by 9 µm silica dust. We found that the prominent silica feature remains stable compared to the data taken 15 years earlier by Spitzer, and that the system lacks large-scale (30%) variation over a 40-year baseline since the IRAS measurement taken in 1983. Characteristics derived from the observed dust mineralogy and SED modeling suggest that the dust has high crystalline mass fraction preferentially in both pyroxene and silica types with the latter mostly confined within sub-au region. The JWST spectrum also reveals surprising hot molecular gas emission from CO2 and CO with traces of other volatiles such as H2O and NH3, apparently stable as seen in the Spitzer data and co-located with the silica dust. We explored several pathways explaining the observed properties of dust and gas, and found that a past giant impact event can simultaneously explain both the large amount of highly processed small grains (crystalline silicates and silica) and volatile emission.
In the preferred giant impact scenario, at least one of the involved bodies needs to have volatile content similar to the carbonaceous chondritic material in the solar system. Volatiles are expected to form as (1) the outgassing products either directly from the newly formed, molten planet or from the large planetesimal remnants of impact ejecta, or (2) the stripped atmospheric material. Tiny silica grains can form either by direct condensation from the impact-produced vapor or by hyper-velocity grinding within pebble-size, vapor condensates. If one of the impacting bodies had a surface ocean before the violent collision, the post ocean impact event also favors the formation of ultrafine particles, which can also provide effective UV shielding for the volatiles.
Alternatively, the tiny silica and molecular gas have no direct relation and are created by non-related processes. Volatiles can be delivered into the inner planetary region from scattered icy bodies originally stored in the outer reservoir. Erosion of a rocky body in a high temperature environment can produce tiny silica dust by degradation of silicates. The keys for these alternatives are whether these processes can be efficient enough to account for the amount of silica and molecular gas observed in the system. Future observations to better determine the gas properties, search for other warm/cold gas and dust, assess the temporal change in both the dust composition and volatiles, and identify the shielding mechanisms against photodissociation will help to shed more light into the EDD phenomenon.
@@ A.1 Stellar Photosphere and Subtraction
For the stellar emission, we used a Kurucz model of =6500 K and logg=4.5, normalized to the optical and JHKs photometry corrected by an interstellar extinction of =0.1 (Román-Zúñiga et al., 2023). Given a distance of 139 pc, the integrated stellar luminosity is 3 L☉ and the stellar radius is 1.4 R☉. These values are all consistent with the values found in the literature for a 1.3 M☉ young star (Rodriguez et al., 2012; Fu et al., 2022; Román-Zúñiga et al., 2023). This stellar model was used for deriving IR excess, computing thermal equilibrium dust temperatures in SED modeling, and photodissociation rate of molecules.
The stellar emission contributes to the majority of the emission shortward of 2.2 µm. Using the Kurucz model, we determine the star has a integrated flux of 153 and 99 mJy for the Spitzer/IRAC I1 and I2 bands, respectively, but 170 and 95 mJy for WISE W1 and W2 bands (due to the different filter profiles). HD 23514 has been monitored by the Kepler and TESS missions. Gaia DR3 catalog gives Teff, logg, [Fe/H], RV as 6319 K, 4.21, -0.24 and 6.320.44 km s-1. Tsantaki et al. (2022) lists its RV as 5.180.28 km s-1 and Mermilliod et al. (2009) lists its V sin as 49.55 km s-1. The system shows no flaring behavior (Ilin et al., 2019), nor Scuti-like pulsations (Bedding et al., 2023) over the monitoring periods. Assuming the star is stable over time, Appendix B discusses the IR variability of the HD 23514 system using archival data from Spitzer and WISE. Figure 6 shows the comparison in the MRS spectrum before and after stellar subtraction, suggesting that the IR excess has a roughly equal contribution as the stellar photosphere at 5 µm (the shortest wavelength of MRS).
A.2 Dust Emissivity & Silica Properties

### v2

@@ 3.4 Is CO2 present in the Spitzer/IRS spectrum?
Inspired by the discovery of CO2 gas, we went back to examine whether a similar gas emission was also present in the Spitzer IRS spectrum taken 15 yr earlier. We retrieved the calibrated spectrum from the Combined Atlas of Sources with Spitzer IRS Spectra (CASSIS)222https://cassis.sirtf.com/atlas/ (Lebouteiller et al., 2011) site. We used the optimal extraction product best for point sources and adopted the uncertainty including both statistical (rms) and systematic errors. The IRS data shown in Figure 3 do show a bump near 15 µm, which was not found in other similar debris systems (for details see Appendix C.4), suggesting that the bump is likely astrophysical in nature. We then determined the general continuum by fitting the CO2 complex region using a polynomial function and forcing the flux between 13.9 and 15.8 µm to be relatively flat for a conservative approach as shown in Figure 3. We then constructed a low-resolution model spectrum by (1) adding the JWST CO2 model (from Section 3.3) to the continuum, and (2) resampling the wavelength grid to match the Spitzer IRS spectral resolution333see the IRS Instrument Handbook for details.. As shown in Figure 3, the low-resolution model spectrum resembles the observed one well if the same level of the CO2 emission detected by JWST was also present. In this case, the strong CO2 Q-branch was detected at 4.3 statistically (2.5 if using the combined error) while the others were buried in the noise. In summary, the same level of the hot CO2 emission is also present in the Spitzer data taken 15 yr earlier and remains unchanged between the two epochs.
@@ 4.1 Colocation of silica and hot gas
The HD 23514 system has been observed with the near-IR adaptive optics imaging at the 10 m Keck II telescope (Rodriguez et al., 2012) and with ALMA at 1.3 mm using a beam size of 15 (Sullivan et al., 2022), and neither resolves nor detects the disk. Despite the high degeneracy, the disk extent can be estimated using spectral energy distribution (SED) models because the bulk of disk emission is sensitive to its temperature (i.e., location for a given dust composition). Vican et al. (2016) characterized the HD 23514 system as a two-temperature disk with a hot component having a dust temperature () of 1082 K and a warm component with of 168 K, roughly consistent with our estimated temperatures for the featureless dust continuum (750 + 200 K; Appendix A). Given the stellar properties, these temperatures correspond to radial locations of 0.1 and 4.8 au, respectively, for blackbody emitters. For imperfect emitters like dust grains, the probable locations are likely to be larger (Figure 9). Under the optically thin condition, the contrast/ratio between the prominent 10 and 20 µm dust features is a good proxy for dust temperature, where the larger the ratio the hotter the dust due to Planck emission (Figure 9). In other words, the bulk of the dust emission that gives rise to the silica feature most likely comes from the hot rather than the warm component.
We explored SED models, aiming to constrain the location of the silica dust (Figure 4, and for details see Appendix D.1). Although such models might be degenerate, they provide diagnostic constraints under simplified assumptions. We found that the minimum grain size in the disk needs to be 0.1–0.5 µm in order to reproduce the 10/20 µm feature strengths, which is smaller than the typical blowout size, 1 µm for HD 23514 (Appendix D.2). Our SED models suggest that the disk is compact (0.1–3 au), and the highly refractory silica grains are mostly confined within the sub-au region, where the temperature is close to their sublimation temperature, in order to reproduce the prominent 9 µm feature and 3–5 µm flux. The system’s IR fractional luminosity is 1.76 with 70% coming from the featureless dust continuum. The total dust mass is g (1.67) and the feature-producing grains account for 5% of the dust mass with only 1% coming from tiny silica dust. Based on the model parameters for the gas (Section C.1) and the silica dust, they are likely colocated very close to the star at sub-au distance.
@@ 4.2 Origins of the Stable Silica Dust
The origin of the large amount of submicron grains in debris disks has been long debated in the literature. The typical radiation-pressure blowout size is 1 µm around HD 23514 except for highly porous grains (Arnold et al., 2019), i.e., the grains that give rise to the prominent mid-IR feature have very short lifetimes (less than a year, Appendix D.3). This implies a very high mass-loss rate (51022 g yr-1) if there were no optical depth effects either to prevent them leaving the system by shielding or to replenish them fast enough (e.g., Thebault & Kral 2019). If this high rate was sustained through collisional cascades over the age of the system (150 Myr), HD 23514 would need to be born with a disk that is 35 times of the minimum mass of solar nebula (MMSN), which is not impossible but unlikely. Instead, the large amount of small grains is more likely to be created by transient events where the high rate is only applicable for a limited recent time. The maximum dust production rates in solar system comets can reach 1013 g yr-1 when they are very close to the Sun. It is then unlikely to sustain the high dust production/loss rate by inward-scattered comets if these grains are continuously being ejected by radiation pressure.
@@
One easy way to address this dilemma is that all or some portion of the small grains are not subject to radiation-pressure blowout, which can be achieved by certain grain properties or being trapped by other means (Rieke et al., 2016; Lebreton et al., 2013; Pearce et al., 2020). The prominent 9 µm feature is thought to come from obsidian or annealed/fused quartz grains (Fujiwara et al., 2012) whose properties make them transparent to stellar photons, therefore not subject to blowout, when they are tiny enough (Artymowicz, 1988). Giant impacts provide perfect conditions for the tiny grain formation. These tiny grains could have been generated either by direct condensation from impact-produced vapor (Johnson & Melosh, 2012) or by hypervelocity grinding within optically thick clouds of debris produced by violent collisions of large bodies (Johnson et al., 2012). In the former case, these silica grains were born tiny; while in the later case, tiny grains were formed through collisional cascades and the optical depth effects within the formation region allow grains that are much smaller than the blowout size to form. As a result, the solid-state feature rising from the tiny silica dust would remain stable until they slowly drift inward due to the Poynting-Robertson drag (Su et al., 2020). The feature would gradually weaken due to dust sublimation once the grains reach the sublimation temperature under timescales of a few hundred years in HD 23514.
@@
In addition, the giant-impact ejecta would also form a population of planetesimals, maintaining a high level of dustiness and creating temporal variation (Watt et al., 2024) as probed by 3–5 µm photometry. The observed variability levels require a minimal corresponding area change in the dust cross section to be on the order of (1–2.5)10-3 au2 (for details see Appendix B). This dust cross section is similar to the emitting area of hot gas, pointing a potential link between the dust and molecular gas production in the sub-au region. Assuming the molecular gas and small dust share the same sub-au region, the estimated gas and dust masses suggest that the small grains could have low (1) Stokes number and coupled with the gas (highly depending on the estimated CO mass, see Appendix C.1). The presence of molecular gas would increase the dust-clearing timescale due to the combination of radiation pressure and gas drag (Takeuchi & Artymowicz, 2001; Kenyon et al., 2016). Exploring grain trapping is beyond the scope of this paper, but it is likely to be an important component to a comprehensive model for the HD 23514 system, particularly in light of the new discovery of gas in the system.
@@ 4.3 Origins of the Volatile Material
In situ formation In the context of the giant-impact hypothesis, molecular gas could be produced by outgassing processes if one or both the impacting bodies are volatile rich. Earth’s early atmosphere was thought to form by retention of these outgassing products (Abe & Matsui, 1985). Impact experiments and model calculations both show that H2O and CO2 are the major gases produced by outgassing of carbonaceous chondritic material over a wide range of temperatures and pressures (e.g., Lange & Ahrens 1982; Schaefer & Fegley 2010; Thompson et al. 2021, and references within). Furthermore, shock-recovery experiments targeting a mixture of olivine, iron, and water also reveal preferential formation of ultrafine (tens of nanometers) smoke particles of silicates and metal oxides, facilitated by the role of supercritical water (Furukawa et al., 2007). These tiny grains might be ubiquitous in post-ocean-impact events and provide effective UV shielding for volatiles. Alternatively, volatiles could also be released from (1) an atmospheric stripping event as has been suggested for HD 172555 (Schneiderman et al., 2021), or (2) large remnants of a giant-impact ejecta as long as they are large enough to retain a significant fraction of volatiles. It is then interesting to note that the change in the dust cross section (derived in Appendix B) is on the same order as the emitting area of the molecular gas, indicating that collisional activities in the aftermath of giant-impact events could be tied to the formation/protection of molecular gas in the inner region.
@@
Delivery from outer region This delivery scenario has been widely accepted in both our solar system (comets) and other debris systems (exocomets) – scattered icy bodies from the outer region releasing volatiles and dust grains as they warm up on their inward migration. In this case, the detected molecular gas might have no direct relationship with the giant-impact-produced small grains observed in the system. For HD 23514, the inward delivery rate might be much lower than that of the debris systems that host Kuiper-belt analogs such as Pic (Beust et al., 1996; Matrà et al., 2019) and Corvi (Marino et al., 2017) because it has a limited volume of cold reservoir for storing icy bodies (lack of a massive Kuiper-belt analog, see Appendix D). This scenario also needs a perturber either internally or externally to efficiently scatter icy bodies into the inner region. Ground-based high-contrast imaging found no other sources within 10″ around HD 23514 except for the M7 brown dwarf companion (Rodriguez et al., 2012) 264 away, while a recent radial velocity (RV) study rules out massive giant planets within 0.1 au (Takarada et al., 2020). External companions under some favorable conditions can excite the planetesimal population via the eccentric Kozai–Lidov mechanism, leading to inner dust and gas production (Naoz, 2016; Young & Wyatt, 2024). Without detailed orbital parameters, it is not clear whether the M dwarf companion at 360 au away is capable of this role. Other challenges in the delivery scenario are: (1) evaporation of icy bodies is expected to occur at a range of distances, which is at odds with the gas concentration within the sub-au region, and (2) the comet delivery rate is expected to be variable based on the solar system observations, which somehow contradicts the tentative CO2 detection in the Spitzer data. The second challenge might be accommodated if scattering of many comets occurs at high efficiency. A future investigation is needed to further explore the delivery model; nonetheless, it remains a possible origin for the molecular gas.
@@ 5 Conclusion
We have presented a JWST MIRI/MRS 5–28 µm spectrum of HD 23514, an F5V star in the 150 Myr-old Pleiades cluster, hosting one of the giant-impact disks discovered by Spitzer with unique mineralogy dominated by 9 µm silica dust. We found that the prominent silica feature remains stable compared to the data taken 15 yr earlier by Spitzer, and that the system lacks large-scale (30%) variation over a 40-year baseline since the IRAS measurement taken in 1983. Characteristics derived from the observed dust mineralogy and SED modeling suggest that the dust has high crystalline mass fraction preferentially in both pyroxene and silica types with the latter mostly confined within sub-au region. The JWST spectrum also reveals surprising hot molecular gas emission from CO2 and CO with traces of other volatiles such as H2O and NH3, apparently stable as seen in the Spitzer data and colocated with the silica dust. We explored several pathways explaining the observed properties of dust and gas, and found that a past giant-impact event can simultaneously explain both the large amount of highly processed small grains (crystalline silicates and silica) and volatile emission.
In the preferred giant-impact scenario, at least one of the involved bodies needs to have volatile content similar to the carbonaceous chondritic material in the solar system. Volatiles are expected to form as (1) the outgassing products either directly from the newly formed, molten planet or from the large planetesimal remnants of impact ejecta, or (2) the stripped atmospheric material. Tiny silica grains can form either by direct condensation from the impact-produced vapor or by hypervelocity grinding within pebble-size vapor condensates. If one of the impacting bodies had a surface ocean before the violent collision, the post-ocean-impact event also favors the formation of ultrafine particles, which can also provide effective UV shielding for the volatiles.
Alternatively, the tiny silica and molecular gas have no direct relation and are created by nonrelated processes. Volatiles can be delivered into the inner planetary region from scattered icy bodies originally stored in the outer reservoir. Erosion of a rocky body in a high temperature environment can produce tiny silica dust via the degradation of silicates. The keys for these alternatives are whether these processes can be efficient enough to account for the amount of silica and molecular gas observed in the system. Future observations to better determine the gas properties, search for other warm/cold gas and dust, assess the temporal change in both the dust composition and volatiles, and identify the shielding mechanisms against photodissociation will help to shed more light into the EDD phenomenon.
@@ A.1 Stellar Photosphere and Subtraction
For the stellar emission, we used a Kurucz model of =6500 K and logg=4.5, normalized to the optical and photometry corrected by an interstellar extinction of =0.1 (Román-Zúñiga et al., 2023). Given a distance of 139 pc, the integrated stellar luminosity is 3 and the stellar radius is 1.4 . These values are all consistent with the values found in the literature for a 1.3 young star (Rodriguez et al., 2012; Fu et al., 2022; Román-Zúñiga et al., 2023). This stellar model was used for deriving IR excess, computing thermal equilibrium dust temperatures in SED modeling, and photodissociation rates of molecules.
The stellar emission contributes to the majority of the emission shortward of 2.2 µm. Using the Kurucz model, we determine the star has an integrated flux of 153 and 99 mJy for the Spitzer/IRAC I1 and I2 bands, respectively, but 170 and 95 mJy for WISE W1 and W2 bands (due to the different filter profiles). HD 23514 has been monitored by the Kepler and TESS missions. Gaia DR3 catalog gives , log , [Fe/H], RV as 6319 K, 4.21, -0.24 and 6.320.44 km s-1. Tsantaki et al. (2022) lists its RV as 5.180.28 km s-1 and Mermilliod et al. (2009) lists its sin as 49.55 km s-1. The system shows no flaring behavior (Ilin et al., 2019), nor Scuti-like pulsations (Bedding et al., 2023) over the monitoring periods. Assuming the star is stable over time, Appendix B discusses the IR variability of the HD 23514 system using archival data from Spitzer and WISE. Figure 6 shows a comparison in the MRS spectrum before and after stellar subtraction, suggesting that the IR excess has a roughly equal contribution as the stellar photosphere at 5 µm (the shortest wavelength of MRS).
A.2 Dust Emissivity and Silica Properties